OpenStudy (jerudavis):

Martha is covering kitchen shelves with shelving paper. She has 6 shelves that are each 1 3/4 feet long. She buys 13 1/4 feet of shelving paper of the correct width. Which equation can be used to determine how much paper she will have left over?

OpenStudy (jerudavis):

@pizza12

OpenStudy (jerudavis):

@Shalayla4

OpenStudy (jerudavis):

@They_Call_Me_Narii

OpenStudy (littleheron):

@hakunamatata34

OpenStudy (jerudavis):

@WildQueen

OpenStudy (jerudavis):

A) 6(13 1/4) = S B) 13 1/4 + 1 3/4 - 6 = S C) 13 1/4 - 6(1 3/4) = S D) 13 1/4(6) + 1 3/4 = S

OpenStudy (jerudavis):